Hi
today nigel is twice as old as jack.4 year s ago, nigel was 3 times as old as jack. How old will nigel
be in another 5 years.
thanks

Answer by jorel555(1290)

Let j be jack. Then nigel is 2j. So
2j-4=3(j-4)
2j-4=3j-12
j=8
jack is now 8; so nigel is 16
16+5=21, which is how old nigel will be in another 5 years. ☺☺☺☺